# Lens: A Foundation Model for Network Traffic

### v1

@@ 4.3 Ablation Studies
In this study, we investigate the influence of various components on the performance of our proposed foundation model in traffic understanding and generation tasks using the USTC-TFC2016 dataset. More specifically, we examine the effects of pre-training, Masked Span Prediction (MSP), Packet Order Prediction (POP), and Homologous Traffic Prediction (HTP) on model’s .
@@
     w/o Pretraining 0.9799 0.9798
Table 7: Ablation studies on understanding task. It can be observed that pre-training with all designed tasks contribute to performance improvement.
@@ 5 Conclusion and Future Work
In this paper, we introduced Lens, a foundation model for network traffic excelling in both network traffic understanding and generation. Built on the T5 architecture, Lenscan effectively learn the latent representations by pre-training on a large-scale unlabeled data, and then fine-tunes on a small amount of data for various downstream tasks. We also explored three different tokenization algorithms on the pre-training performance, and found that WordPiece with a predefined vocabulary performed the best. To further enhance the performance of Lens, we designed a new objective function that considers three pre-training tasks. Besides, we did ablation studies to validate the effectiveness of proposed components in the pre-training process. We evaluated the performance of Lenson 6 real-world datasets across 15 understanding tasks and 5 header fields generation tasks. Evaluation results demonstrated that our model outperforms the baselines in most of downstream tasks, showcasing its effectiveness and strong generalization ability. It is worth noting that our approach required 50% $\sim$ 95% less labeled data for fine tuning compared to other pre-training methods like ET-BERT and YaTC.
For future work, we plan to increase the input window size and upscale the model parameters for Lens. What is more, we would like to investigate trustworthy foundation models for network traffic.
@@ A.1.2 Performance Analysis
As shown in Table 9 and 10, we can see that (1) Lensoutperforms the NetShare according JSD and TVD metrics by a large margin in most of tasks, indicating that the distribution of results generated by Lensis closer to the real settings than that generated by NetShare; (2) Lenscan generate more diversified IP addresses compared to NetShare, meaning that Lenscan mimic the real-world network complexity better.
JSD TVD DR JSD TVD DR JSD TVD DR JSD TVD DR JSD TVD DR JSD TVD DR JSD TVD DR
Table 9: Comparison results of source IP generation task.
JSD TVD DR JSD TVD DR JSD TVD DR JSD TVD DR JSD TVD DR JSD TVD DR JSD TVD DR

### v3

@@ 4.3 Ablation Studies
In this study, we investigate the influence of various components on the performance of our proposed foundation model in traffic understanding and generation tasks using the USTC-TFC2016 dataset. More specifically, we examine the effects of pre-training, Masked Span Prediction (MSP), Packet Order Prediction (POP), and Homologous Traffic Prediction (HTP) on model’s performance .
@@
Table 7: Ablation studies on understanding task. It can be seen that pre-training with all designed tasks contribute to performance improvement.
@@
       w/o Pre-training 0.9799 0.9798
@@ 5 Conclusion and Future Work
In this paper, we introduced Lens, a foundation model for network traffic excelling in both network traffic understanding and generation. Built on the T5 architecture, Lens  can effectively learn the latent representations by pre-training on large-scale unlabeled data. We also explored three different tokenization algorithms on the pre-training performance, and found that WordPiece with a predefined vocabulary performs the best. To further enhance the performance of Lens, we designed a new objective function that takes three pre-training tasks into account. Besides, we conducted ablation studies to validate the effectiveness of proposed components in the pre-training process. We evaluated the performance of Lens  on 6 real-world datasets, including 15 understanding tasks and 5 header fields generation tasks. Evaluation results demonstrated that our model outperforms the baselines in most of downstream tasks, showcasing its effectiveness and strong generalization ability. It is worth noting that our approach required 50% $\sim$ 95% less labeled data for fine-tuning compared to other pre-training methods like ET-BERT and YaTC.
For future work, we plan to upscale the model parameters for Lens. Moreover, we would like to investigate trustworthy foundation models for network traffic.
@@ A.1.2 Performance Analysis
As shown in Table 10 and 11, we can see that (1) Lens  outperforms the NetShare according JSD and TVD metrics by a large margin in most of tasks, indicating that the distribution of results generated by Lens  is closer to the real settings than that generated by NetShare; (2) Lens  can generate more diversified IP addresses compared to NetShare, meaning that Lens  can mimic the real-world network complexity better.
@@ A.2 Visualizing Traffic Generation Distribution Results
A.3 Experiments over Pre-training Loss Hyperparameters
To determine the best weight $\alpha$ and $\beta$ in Eq. (4), we experiment some hyperparameter settings as shown in table 9. We can see from this table that when $\alpha$ and $\beta$ are all set to $0.2$, Lens  achieves the best performance on Masked Span Prediction task.
Table 9: Accuracy of Masked Span Prediction under different $\alpha$ and $\beta$ settings in equation 4.
$\beta$ $\alpha$ 0.1 0.2 0.5 1.0
0.1 0.654 0.431 0.649 -
0.2 0.649 0.689 0.634 -
0.5 0.596 0.655 0.599 -
1.0 - - - 0.462
@@ A.3 Experiments over Pre-training Loss Hyperparameters
Table 10: Comparison results of source IP generation task.
JSD TVD DR JSD TVD DR JSD TVD DR JSD TVD DR JSD TVD DR JSD TVD DR JSD TVD DR
JSD TVD DR JSD TVD DR JSD TVD DR JSD TVD DR JSD TVD DR JSD TVD DR JSD TVD DR